# When is the canonical conductor minimal?

### v1

@@ Abstract.
For a one dimensional analytically unramified local domain $R$, the blowup algebra of the canonical ideal is a birational extension. The conductor of this extension always contains the conductor of $R$. We study the case when there is equality and we classify numerical semigroups with this property.
@@ 1. Introduction
In this paper, we are treating the missing case: the case where the equality $b(\omega)=\mathrm{co}(R)$ holds. We say that such a ring has minimal canonical conductor. By definition, the class of such rings contain the class of far-flung Gorenstein rings. Under some conditions, they coincide. For example, if $R$ has multiplicity 3, then $R$ has minimal canonical conductor if and only if it is far-flung Gorenstein. But in general the class of rings with minimal canonical conductor is much larger. In Section 3, we provide an example of a ring with minimal canonical conductor which is nearly Gorenstein.
In Section 2, we recall some preliminaries and prove equivalent conditions for having minimal conductor for an analytically unramified one dimensional local domain. Our main result in this section is the following.
Theorem A (Theorem 2.11).
Let $R$ be a one dimensional singular local domain which is analytically unramified. Let $\omega$ be a canonical ideal. Then, the following are equivalent.
  1. (1)
    The blowup algebra $B(\omega)$ is the normalisation $\overline{R}$.
  2. (2)
    There is an equality $b(\omega)=\mathrm{co}(R)$.
  3. (3)
    For any $\omega$-Ulrich module $M$, we have an isomorphism $M\cong\overline{R}^{\oplus r}$ for some $r$.
  4. (4)
    There exists an $n$ such that $\omega^{n}\cong\mathrm{co}(R)$.
In Section 3, we apply this theorem to numerical semigroup rings and classify numerical semigroup rings with minimal multiplicity.
Theorem B.
Let $H$ be a numerical semigroup with Frobenius number $f$. Then, the ring $k[[H]]$ has minimal canonical conductor if and only if f1f-1 is a quasi-Frobenius number. In particular, if $H$ is minimally generated by $n<a_{2}<\ldots<a_{n}$, then the ring $k[[H]]$ has minimal canonical conductor if and only if $a_{n}=a_{n-1}+1$.
@@ Acknowledgements
I would like to thank Alfred Geroldinger, Alessio Moscariello and Shinya Kumashiro for valuable conversations and helpful comments.
2. Preliminaries and the main theorem
We start this section with some conventions and general definitions. Even though we give some definitions in a larger generality, we will always work with one dimensional Noetherian local domains throughout this paper and all modules will be finitely generated. Hence, our rings will be Cohen-Macaulay local rings and they will be generically Gorenstein. So, they will possess a canonical module.
We start with recalling some definitions.
@@ 2. Preliminaries and the main theorem
Let $R$ be a one dimensional local domain which is analytically unramified. Assume that a finitely generated $R$-module has a nonzero trace ideal $\operatorname{\mathrm{tr}}(M)$ with a principal reduction. Then, there exists an inclusion $\mathrm{co}(R)\subseteq\operatorname{\mathrm{tr}}(M)$. When $R$ has infinite residue field, the principal reduction condition is satisfied automatically.
The following lemma was established by Faber [Fab20, Theorem 3.5] in the reduced case and more generally by Dao-Maitra-Sridhar [DMS23, Theorem 2.9].
Lemma 2.4.
Let $R$ be a commutative Noetherian ring, $S$ a finite birational extension and $M$ a finitely generated $R$-module. Consider the two statements:
  1. (1)
    $M$ is a module over $S$.
  2. (2)
    There is an inclusion $\operatorname{\mathrm{tr}}(M)\subseteq\mathrm{co}(R,S)$.
Then, (1) implies (2). When $M$ is reflexive, (2) implies (1), as well.
This gives us the following lemma about reflexive modules with minimal traces.
Lemma 2.5.
Let $R$ be a one dimensional local domain and assume that it is analytically unramified and $M$ a reflexive $R$-module of constant rank $r>0$. If there exists an equality $\operatorname{\mathrm{tr}}(M)=\mathrm{co}(R)$, then we have an isomorphism
$M\cong\overline{R}^{\oplus r}.$
Proof.
Let $M$ be a reflexive $R$-module with $\operatorname{\mathrm{tr}}(M)=\mathrm{co}(R)$. Then by Lemma 2.4, we know that $M$ is a module over $\overline{R}$. Since $M$ is reflexive, it is a maximal Cohen-Macaulay $R$-module and hence a maximal Cohen-Macaulay $\overline{R}$-module. Since $R$ is a one dimensional local domain, we know that $\overline{R}$ is a regular local ring. Hence, any maximal Cohen-Macaulay module over $\overline{R}$ is a direct sum of copies of $\overline{R}$. This finishes the proof. ∎
@@
Let $R$ be a one dimensional Cohen-Macaulay local ring and $I$ a regular ideal. Then a maximal Cohen-Macaulay $R$-module is called $I$-Ulrich if there is an isomorphism $IM\cong M$.
@@
We are now ready to state and prove the main theorem of this section.
@@
Let $R$ be a one dimensional singular local domain which is analytically unramified. Let $\omega$ be a canonical ideal. Then, the following are equivalent.
@@
    For any $\omega$-Ulrich module $M$, we have an isomorphism $M\cong\overline{R}^{\oplus r}$ for some $r$.
@@
    There exists an $n$ such that $\omega^{n}\cong\mathrm{co}(R)$.
Proof.
If two modules are isomorphic, then they have the same trace ideal which shows the implication $(1)\implies(2)$. If $M$ is an $\omega$-Ulrich module, then we have
$\mathrm{co}(R)\subseteq\operatorname{\mathrm{tr}}(M)\subseteq b(\omega)=\mathrm{co}(R).$
With this, Lemma 2.5 shows the implication $(2)\implies(3)$. Now assume $(3)$. By Remark 2.10(3), for $n\gg 0$, we know that $\omega^{n}$ is $\omega$-Ulrich. Hence, $\omega^{n}\cong\overline{R}^{\oplus r}\cong\mathrm{co}(R)^{\oplus r}$ for some $r$. As $\omega^{n}$ has rank 1, $r$ must be $1$ which proves $(4)$. Now, assume $(4)$. Then, we have equalities $B(\omega)=B(\omega^{n})=B(\mathrm{co}(R))=\overline{R}$. This implies $(1)$ showing that all conditions are equivalent. ∎
We end this section with our definition and a remark on relationship between far-flung Gorenstein rings.
@@
Let $R$ be a one dimensional singular local domain with canonical ideal $\omega$. Assume that $R$ is analytically unramified. We say that $R$ has minimal canonical conductor if the equivalent conditions of Theorem 2.11 hold for $R$.
@@
Let $n\geq 3$ be a positive integer and $R=k[[t^{n},t^{n+1},\ldots,t^{2n-1}]]$. Then, one has equalities
$\mathrm{co}(R)=b(\omega)=\operatorname{\mathrm{tr}}(\omega)=\mathfrak{m}$
and $R$ has minimal canonical conductor and the canonical conductor is equal to the maximal ideal. Rings with this property have been considered by several authors in the last decade. For instance, they appear in Faber’s work [Fab20] as one-step normal rings. In the work of Dao-Dey-Dutta [DDD22], such rings are exactly one dimensional Ulrich split rings. In the language of Herzog-Kumashiro-Stamate [HKS23], they are far-flung Gorenstein and nearly Gorenstein. Recently, such rings also appeared in the work of Geroldinger-Yan-Zhang as conductor domains [GYQ].
Remark 2.14.
As we always have the inclusions $\mathrm{co}(R)\subseteq b(\omega)\subseteq\operatorname{\mathrm{tr}}(\omega)$ for a one dimensional analytically unramified local domain, the condition $\operatorname{\mathrm{tr}}(\omega)=\mathrm{co}(R)$ implies the condition $b(\omega)=\mathrm{co}(R)$. Hence, far-flung Gorenstein rings have minimal canonical conductor. However, not all rings with minimal canonical conductor are far-flung Gorenstein. We will see examples in the next section. On the other hand, we have the next proposition which shows when the two classes agree.
Before stating the proposition, we need a definition due to Kumashiro [Kum22].
@@
For a one dimensional local domain, the following are equivalent.
@@
The inclusion $b(\omega)\subseteq\operatorname{\mathrm{tr}}(\omega)$ is an equality if and only if there exists an isomorphism $\operatorname{\mathrm{tr}}(\omega)\cong\omega^{*}$ by [DL24, Proposition 4.7]. And such an isomorphism holds if and only if $R$ has canonical reduction number at most 2 by [Kum22, Theorem 2.13]. This shows that if $R$ has minimal canonical conductor and canonical reduction number 2, then it is far-flung Gorenstein. On the other hand, as mentioned above all far-flung Gorenstein rings have minimal canonical conductor. Moreover, by [HKS23, Theorem 2.5], we see that $R$ has canonical reduction number 2. This finishes the proof. ∎
We finish this section with an observation on the endomorphism ring $\mathfrak{m}\colon\mathfrak{m}$. This ring is often useful when one studies reflexive modules over $R$ and is far-flung Gorenstein when $R$. The latter assertion was proved by Herzog-Kumashiro-Stamate in [HKS23, Theorem 3.3] and the proof of the proposition below works similarly by making use of Claim 1 in the proof of op. cit.
@@
Let $(R,\mathfrak{m})$ be a one dimensional local domain. If $R$ has minimal canonical conductor, then the ring $B=\mathfrak{m}\colon\mathfrak{m}$ also has minimal canonical conductor.

### v2

@@ Abstract.
For a one dimensional analytically unramified Cohen-Macaulay local ring $R$, the blowup algebra of the canonical ideal is a module finite birational extension. The conductor of this extension always contains the conductor of $R$. We study the case when there is equality. This is the case where $R$ is far from being almost Gorenstein. We study this property within the landscape of numerical semigroup rings and local Arf rings.
@@ 1. Introduction
In this paper, we are treating the missing case: the case where the equality $b(\omega)=\mathrm{co}(R)$ holds. We say that such a ring has minimal canonical conductor. By definition, the class of such rings contain the class of far-flung Gorenstein rings. Under some conditions, they coincide. For example, if $R$ has multiplicity 3, then $R$ has minimal canonical conductor if and only if it is far-flung Gorenstein. But in general the class of rings with minimal canonical conductor is much larger.
In Section 2, we recall some preliminaries and equivalent conditions which we use to define a ring with minimal canonical conductor. In Section 3, we compare this property with being far-flung Gorenstein and also study this property for blowups at the maximal ideal.
Section 4 compares having minimal canonical conductor with having reflexive and Gorenstein birational extensions. We then apply this to understand local Arf domains which have minimal canonical conductor.
Finally, in Section 5 we classify numerical semigroup rings which have minimal canonical conductor and we give an example which has minimal canonical conductor but is nearly Gorenstein.
@@ Acknowledgements
I would like to thank Alfred Geroldinger, Alessio Moscariello and Shinya Kumashiro for valuable conversations and helpful comments. I also want to thank Souvik Dey for his comments after the first preprint version of this paper.
2. Preliminaries
We start this section with some conventions and general definitions.
@@ 2. Preliminaries
    Here, by a birational extension we mean a ring extension $R\subseteq S\subseteq Q(R)$.
@@
Recall that a local ring is called analytically unramified if its completion is reduced. For an analytically unramified local ring, the normalisation $\overline{R}$ is finitely generated as an $R$-module. The converse is also true when $R$ is a one dimensional Cohen-Macaulay local ring. In this case, $R$ is reduced.
@@
Let $R$ be a one dimensional Cohen-Macaulay local ring which is analytically unramified. Assume that for a finitely generated $R$-module M, the trace ideal $\operatorname{\mathrm{tr}}(M)$ contains a nonzerodivisor and has a principal reduction. Then, there exists an inclusion $\mathrm{co}(R)\subseteq\operatorname{\mathrm{tr}}(M)$ [DMS23, Corollary 3.6]. When $R$ has infinite residue field, the principal reduction condition is satisfied automatically.
@@
Let $R$ be a one dimensional Cohen-Macaulay local ring and $I$ a regular ideal. Then a maximal Cohen-Macaulay $R$-module is called $I$-Ulrich if there is an isomorphism $IM\cong M$. We denote by $\mathrm{Ul}_{I}(R)$ the category of $I$-Ulrich modules.
@@
The following theorem follows from [DDD22, Theorem 5.5]. We note that by the additive closure $\mathrm{add}_{R}(X)$ of an $R$-module $X$, we mean the smallest subcategory that contains $X$ and is closed under direct summands, isomorphisms and finite direct sums.
@@
Let $R$ be a one dimensional Cohen-Macaulay local ring which is analytically unramified. Let $\omega$ be a canonical ideal. Then, the following are equivalent.
@@
    There exists an $n$ such that $\omega^{n}\cong\mathrm{co}(R)$.
These equivalent conditions imply the following.
@@
    There is an equality $\mathrm{Ul}_{\omega}(R)=\mathrm{add}_{R}(\overline{R})$.
When $R$ is a Henselian domain, all four conditions are equivalent.
We end this section with our definition and an example.
@@
Let $R$ be a one dimensional Cohen-Macaulay local ring with canonical ideal $\omega$. Assume that $R$ is analytically unramified and is not regular. We say that $R$ has minimal canonical conductor if the equivalent conditions of Theorem 2.9 hold for $R$.
@@
Let $n\geq 3$ be a positive integer and $R=k[[t^{n},t^{n+1},\ldots,t^{2n-1}]]$. Then, by direct computation we can see that $\mathrm{co}(R)=\mathfrak{m}$ and therefore one has equalities
$\mathrm{co}(R)=b(\omega)=\operatorname{\mathrm{tr}}(\omega)=\mathfrak{m}.$
Therefore, $R$ has minimal canonical conductor. Rings with maximal conductor have been considered by several authors in the last decade. For instance, they appear in Faber’s work [FAB20] as one-step normal rings. In the work of Dao-Dey-Dutta [DDD22], such rings are exactly one dimensional Ulrich split rings. In the language of Herzog-Kumashiro-Stamate [HKS23], they are far-flung Gorenstein and nearly Gorenstein. Recently, such rings also appeared in the work of Geroldinger-Yan-Zhang as conductor domains [9].
3. Minimal canonical conductor versus far-flung Gorenstein
In Example 2.11, we have seen that the canonical conductor $b(\omega)$ equals the canonical trace ideal $\operatorname{\mathrm{tr}}(\omega)$. As we always have the inclusions $\mathrm{co}(R)\subseteq b(\omega)\subseteq\operatorname{\mathrm{tr}}(\omega)$ for a one dimensional analytically unramified local ring, the condition $\operatorname{\mathrm{tr}}(\omega)=\mathrm{co}(R)$ implies the condition $b(\omega)=\mathrm{co}(R)$. Hence, far-flung Gorenstein rings have minimal canonical conductor. However, not all rings with minimal canonical conductor are far-flung Gorenstein. We will see examples in the next section. in this section, we will look at the case when the two classes agree. We start with a definition due to Kumashiro [KUM22].
@@ 3. Minimal canonical conductor versus far-flung Gorenstein
As mentioned above, a one dimensional analytically unramified Cohen-Macaulay local ring is reduced and therefore it is generically Gorenstein.
@@
Let $R$ be a one dimensional Cohen-Macaulay local ring which is analytically unramified. The following are equivalent.
@@
The inclusion $b(\omega)\subseteq\operatorname{\mathrm{tr}}(\omega)$ is an equality if and only if there exists an isomorphism $\operatorname{\mathrm{tr}}(\omega)\cong\omega^{*}$ by [DL24, Proposition 4.7]. And such an isomorphism holds if and only if $R$ has canonical reduction number at most 2 by [KUM22, Theorem 2.13]. This shows that if $R$ has minimal canonical conductor and canonical reduction number 2, then it is far-flung Gorenstein. On the other hand, as mentioned above all far-flung Gorenstein rings have minimal canonical conductor. Moreover, by [HKS23, Theorem 2.5], we see that $R$ has canonical reduction number 2 when it is far-flung Gorenstein. This finishes the proof. ∎
We finish this section with an observation on the endomorphism ring $\mathfrak{m}\colon\mathfrak{m}$. This ring is often useful when one studies reflexive modules over $R$ and is far-flung Gorenstein whenever $R$ is. The latter assertion was proved by Herzog-Kumashiro-Stamate in [HKS23, Theorem 3.3] and the proof of the proposition below works similarly by making use of Claim 1 in the proof of op. cit.
@@
Let $(R,\mathfrak{m})$ be a one dimensional analytically unramified local Henselian domain. If $R$ has minimal canonical conductor, then the ring $B=\mathfrak{m}\colon\mathfrak{m}$ also has minimal canonical conductor.
4. Reflexive and Gorenstein birational extensions
Our definition of a ring with minimal canonical conductor via the chain of ideals
suggests that such a ring should be considered as far away as possible from being almost Gorenstein. However, this claim uses a specific characterisation of one dimensional almost Gorenstein rings, namely [DMS23, Proposition 7.1]. However, One of the most well-known characterisations of a one dimensional almost Gorenstein local ring $(R,\mathfrak{m})$ is via the endomorphism ring $E=\mathfrak{m}\colon\mathfrak{m}\cong\mathrm{End}_{R}(\mathfrak{m})$. Precisely, in [GMP13, Theorem 5.1], Goto-Matsuoka-Phuong proves that the following are equivalent for a one dimensional Cohen-Macaulay local ring $(R,\mathfrak{m})$:
  1. (1)
    $E$ is a Gorenstein ring.
  2. (2)
    $R$ is an almost Gorenstein ring with minimal multiplicity.
In this section, we will explore this point of view. We start with some important remarks.
Remark 4.1.
Let $(R,\mathfrak{m})$ be a one dimensional Cohen-Macaulay local ring which is generically Gorenstein and put $E=\mathrm{End}_{R}(\mathfrak{m})$ as above. Then, both $\mathfrak{m}$ and $E$ are reflexive $R$-modules, see for instance [KOB22, Lemma 2.5]. Assume that $R$ is not regular. Then, by [KOB20, Lemma 2.2], for any module-finite reflexive birational extension $S$ of $R$, one has a chain of inclusions
$R\subsetneq E\subseteq S.$
That is, $E$ is the minimal module-finite reflexive birational extension of $R$, in a sense. Combining this with the fact that any module-finite birational extension of a Gorenstein ring is also Gorenstein, we may rephrase the theorem of Goto-Matsuoka-Phuong as follows: for a one dimensional Cohen-Macaulay local ring $R$, the following are equivalent.
  1. (1)
    Any module-finite reflexive birational extension of $R$ is Gorenstein.
  2. (2)
    $R$ is an almost Gorenstein ring with minimal multiplicity.
Remark 4.2.
Let $(R,\mathfrak{m})$ be a one dimensional Cohen-Macaulay local ring with infinite residue field and assume that it is not regular. In this case, $\mathfrak{m}$ is not a principal ideal and therefore $\mathfrak{m}$ is a regular proper trace ideal, see for instance [HR23, Proposition 2.8]. By [OOI96, Proposition 2.5], we have that $R$ has minimal multiplicity if and only if there exists an isomorphism $\mathfrak{m}\cong\mathfrak{m}^{-1}\cong\mathfrak{m}^{*}$. And by [DL24, Proposition 3.3], the latter isomorphism is equivalent to saying that $\mathfrak{m}$ is a stable trace ideal. Finally, let $\omega$ be the canonical module of $R$ and assume that $R$ is not Gorenstein so that $\omega$ is not free. If $R$ is almost Gorenstein, then by [DMS23, Theorem 7.1], we have $b(\omega)=\mathfrak{m}$. Combining all of these, we reach at the following paraphrase of the Goto-Matsuoka-Phuong theorem. The following are equivalent for a one dimensional Cohen-Macaulay local ring $R$ with canonical module $\omega$:
  1. (1)
    Any module-finite reflexive birational extension of $R$ is Gorenstein.
  2. (2)
    $R$ is an almost Gorenstein ring and $b(\omega)$ is stable.
Let $R$ be a one dimensional analytically unramified local ring. Then, the normalisation $\overline{R}$ of $R$ is a reflexive birational extension which is Gorenstein. Our main theorem in this section proves that this is the only such birational extension when $R$ has minimal canonical conductor.
Theorem 4.3.
Let $R$ be a one dimensional Cohen-Macaulay local ring which is analytically unramified. Consider the two conditions.
  1. (1)
    $R$ has minimal canonical conductor.
  2. (2)
    If $R\subseteq S\subseteq\overline{R}$ is a Gorenstein birational extension that is reflexive as an $R$-module, then $S=\overline{R}$.
Then, (1) implies (2). If $b(\omega)$ is stable, then (2) implies (1).
Proof.
Assume that $R$ has minimal canonical conductor. Let $S$ be a Gorenstein birational extension of $R$ which is reflexive as an $R$-module. Then, by [DMS23, Theorem 7.10], we know that $\mathrm{co}(R,S)$ is $\omega$-Ulrich. Therefore, $\mathrm{co}(R,S)$ belongs to $\mathrm{add}_{R}(\mathrm{co}(R))$ and hence $\mathrm{co}(R,S)=\mathrm{co}(R)$ which shows that $S=\overline{R}$.
On the other hand, assume (2) and that $b(\omega)$ is stable. We know that $B(\omega)$ is an $\omega$-Ulrich module. So, it is a reflexive birational extension of $R$. We know that $\operatorname{\mathrm{tr}}(B(\omega))=\mathrm{co}(R,B(\omega))=b(\omega)$ is $\omega$-Ulrich, as well. Moreover, by our assumption, it is also stable. Thus, by [DMS23, Theorem 7.10], we see that $B(\omega)$ is Gorenstein. Hence, by our assumption, we get that $B(\omega)=\overline{R}$ and therefore, we are done by Theorem 2.9. ∎
We finish this section with an application to Arf local rings. The following definition and remarks on Arf local rings can be found in Lipman’s original work on Arf rings [LIP71].
Definition 4.4.
Let $R$ be a one dimensional Noetherian semi-local Cohen-Macaulay ring.
  1. (1)
    We define a chain of blowup algebras inductively by letting $R_{0}=R$ and $J_{0}=J(R)$ (where $J(R)$ denotes the Jacobson radical of $R$) and setting $R_{n+1}:=\mathrm{End}_{R_{n}}(J_{n})$ where $J_{n}=J(R_{n})$. If $\mathfrak{m}$ is a maximal ideal of some $R_{n}$, then the local ring $(R_{n})_{\mathfrak{m}}$ is said to be infinitely near to $R$.
  2. (2)
    We say that $R$ is an Arf ring if every integrally closed ideal containing a nonzerodivisor has a principal reduction and whenever $x\in R$ is nonzerodivisor and $y,z\in R$ arbitrary elements with $y/x$ and $z/x$ in $\overline{R}$, we must have $yz/x\in R$.
Remark 4.5.
The following are equivalent:
  1. (1)
    $R$ is an Arf ring,
  2. (2)
    Every integrally closed ideal containing a nonzerodivisor is stable,
  3. (3)
    If $B$ is any local ring infinitely near to $R$, then $B$ has minimal multiplicity.
It is clear from this that if $R=R_{0}$ is Arf, then $R_{n}$ is Arf for any $n\geq 0$. Moreover, if $R$ is a local ring with minimal multiplicity and $R_{1}$ is Arf, then $R$ is also Arf.
Let $R$ be a complete local domain of dimension one and assume that it is analytically unramified. Then, we have a chain
$R=R_{0}\subseteq R_{1}\subseteq R_{2}\subseteq\cdots\subseteq R_{n-1}\subseteq R_{n}=\overline{R}$
consisting of complete local domains which are infinitely near to $R$. By [DAO21, Theorem A], the $R_{i}$’s give a complete list of indecomposable reflexive $R$-modules up to isomorphism when $R$ is Arf. Now, assume that $R$ is Arf. If $R\subseteq S\subseteq\overline{R}$ is a birational extension that is reflexive as an $R$-module, then $S=R_{i}$ for some $i=1,\ldots,n$. Hence, any birational extension $S\subseteq\overline{R}$ which is reflexive as an $R$-module is Arf. On the other hand, by [BDF00, Corollary 4.3], $S$ is Gorenstein and Arf if and only if $S$ has multiplicity 2. This gives us the following characterisation by applying Theorem 4.3.
Theorem 4.6.
Let $R$ be a complete local Arf domain and assume that it is analytically unramified. The following are equivalent.
  1. (1)
    $R$ has minimal canonical conductor.
  2. (2)
    For any birational extension $R\subseteq S\subseteq\overline{R}$ that is reflexive as an $R$-module, we have either $S=\overline{R}$ or $e(S)\geq 3$.
  3. (3)
    There does not exist an $A_{n}$-singularity that is infinitely near to $R$.